OpenStudy (anonymous):

evaluate 10!/5!5! A. 1 B. 120 C. 210 D. 252 I think it would be A. 1

OpenStudy (anonymous):

D. 252

OpenStudy (anonymous):

can you explain the steps please

OpenStudy (anonymous):

\[\frac{ 10! }{ 5!*5! }=\frac{ 10*9*8*7*6*5! }{5!*5! }=\frac{ 10*9*8*7*6 }{ 5! }\]\[=\frac{ (2*5)*(3*3)*(2*4)*(7)*(2*3) }{ 1*2*3*4*5 }= (3)*(2)*(7)*(2*3) =42*6=252\]@deigo

OpenStudy (anonymous):

@deigo

OpenStudy (anonymous):

Medal? -.-

OpenStudy (anonymous):

thank you